Lavalear Corporation uses a process costing system and traces costs through several processing departments, starting with the Bonding Department. The following is information taken from the Bonding Department’s September production cost report.

Cost Data: Bonding Department
Direct materials costs in beginning inventory, September 1 $ 50,400
Conversion costs in beginning inventory, September 1 36,000
Direct materials costs incurred in September 789,750
Conversion costs incurred in September 787,500
Physical Units: Bonding Department
Units in process, September 1 300
Units transferred out during September 1,500
Units in process, September 30 500
Percentages of Completion: Bonding Department
Direct materials, September 1 40 %
Conversion, September 1 25
Direct materials, September 30 75 %
Conversion, September 30 30

Exercise 18.14 Part a

a. Compute the cost per equivalent unit of direct materials and conversion carried forward from August and assigned to the Bonding Department's beginning inventory on September 1.
Answer #1

a) Compute the cost per equivalent unit of direct materials and conversion carried forward from August and assigned to the Bonding Department's beginning inventory on September 1.

Equivalent unit of material Beginning inventory = 300*40% = 120 Units

Equivalent unit of conversion beginning inventory = 300*25% = 75 Units

Cost per equivalent unit of direct material = 50400/120 = 420

Cost per equivalent unit of conversion = 36000/75 = 480